# Taming Self-Training for Open-Vocabulary Object Detection

### v2

@@ Abstract
Recent studies have shown promising performance in open-vocabulary object detection (OVD) by utilizing pseudo labels (PLs) from pretrained vision and language models (VLMs). However, teacher-student self-training, a powerful and widely used paradigm to leverage PLs, is rarely explored for OVD. This work identifies two challenges of using self-training in OVD: noisy PLs from VLMs and frequent distribution changes of PLs. To address these challenges, we propose SAS-Det that tames self-training for OVD from two key perspectives. First, we present a split-and-fusion (SAF) head that splits a standard detection into an open-branch and a closed-branch. This design can reduce noisy supervision from pseudo boxes. Moreover, the two branches learn complementary knowledge from different training data, significantly enhancing performance when fused together. Second, in our view, unlike in closed-set tasks, the PL distributions in OVD are solely determined by the teacher model. We introduce a periodic update strategy to decrease the number of updates to the teacher, thereby decreasing the frequency of changes in PL distributions, which stabilizes the training process. Extensive experiments demonstrate SAS-Det is both efficient and effective. SAS-Det outperforms recent models of the same scale by a clear margin and achieves 37.4 AP and 29.1 AP on novel categories of the COCO and LVIS benchmarks, respectively.
@@ 1 Introduction
Traditional closed-set object detectors [6, 16, 39] are restricted to detecting objects with a limited number of categories. Increasing the size of detection vocabularies usually requires heavy human labor to collect annotated data. With the recent advent of strong vision and language models (VLMs) [26, 36], open-vocabulary object detection (OVD) [14] provides an alternative direction to approach this challenge. Typically, OVD detectors are trained with annotations of base categories and expected to generalize to novel target categories by leveraging the power of pretrained VLMs.
One promising thread of recent studies for OVD [11, 12, 57, 49] leverages VLMs to obtain pseudo labels (PLs) beyond base categories. But they rarely explore self-training, a powerful and widely used schema for utilizing PLs in closed-set tasks [46, 45, 51, 52]. We investigate self-training for OVD and find the vanilla self-training approach does not improve OVD performance due to the following challenges.
@@
Second, self-training for closed-set object detection [46, 52, 25] usually follows an online teacher-student manner. In each training iteration, the teacher generates PLs, and the student is trained with a mixture of ground truth and PLs. Then, the teacher is updated by the student with exponential moving average (EMA). However, we find such EMA updates degrade OVD models (see Table 4). Our hypothesis is that, unlike closed-set tasks, OVD provides no ground truth for target categories, and thus, the supervision for target categories is fully decided by the distribution of PLs predicted by the teacher. Hence, the EMA updates change the distribution of PLs in each iteration, making the training unstable.
@@ 2 Related Work
Open-vocabulary object detection (OVD). Traditional closed-set object detectors [38, 39, 16, 6] achieve great performance but are limited in their vocabulary size. Zero-shot methods [3, 37, 60, 53] increase the vocabulary size but with limited accuracy. Motivated by the strong zero-shot abilities of VLMs, recent efforts focus on OVD. Finetuning-based methods [56, 33, 58, 24, 22] add detection heads onto pretrained VLMs and then finetune the detector with concepts of base categories. Such methods are simple but may forget the knowledge learned in the pretraining [2]. Distillation-based methods [14, 10, 48] introduce additional distillation loss functions that force the output of a detector to be close to that of a VLM and thus avoid forgetting. However, since the distillation losses are not designed for the detection task, they may conflict with detection objectives due to gradient conflicts that are a common issue for multi-task models. Methods like Gao et al. [12], Zhao et al. [57], Feng et al. [11], Wu et al. [49] create pseudo labels (PL) of novel concepts as supervision, and do not require extra losses, which sidesteps both catastrophic forgetting and gradient conflicts. But these methods need handcrafted steps to generate high quality PLs, e.g. multiple runs of box regression [57], activation maps [12] from Grad-CAM [42], image retrieval [11], or multi-stage training [49]. In this work, we address the two challenges of using self-training for OVD and enable an efficient end-to-end pseudo labeling pipeline. Besides, we point out the noise due to the poor locations of PLs and introduce the SAF head to handle such noise.
Self-training for object detection. Weakly-supervised object detection methods [46, 47, 40] explore online self-training by distilling the knowledge from the model itself. Recent semi-supervised object detection methods [52, 32, 25] adopt a teacher-student design, where the teacher is an exponential moving average (EMA) of the student. Self-training has been widely explored in the above fields but rarely in OVD. Unlike semi-supervised object detection, OVD encounters two challenges to use self-training, i.e., more noisy PLs, and larger changes in PLs’ distributions. This work proposes the SAF head to address the noise, and adopts periodic updates to reduce the frequency of changes in PLs’ distribution.
@@ 3.1 Adapting CLIP to OVD
In this section, we introduce how to adapt the ResNet-based CLIP into a Faster-RCNN [39] (C4) detector. For simplicity, we do not use FPN [31], but it can be incorporated with learnable gating from Flamingo [1]. Similarly, ViT-based CLIP models can be adapted to detectors following [28].
@@ 3.2 Taming Self-Training
Reducing changes in PLs’ distributions with periodic updates. The teacher model is updated with the student to enable self-training. The exponential moving average (EMA) is a widely used approach for object detection [25, 32, 52], which updates the teacher in every iteration. But we observed empirically that it does not benefit OVD (Table 4). We hypothesize that, unlike semi-supervised object detection, OVD has no ground truth for the target categories. Thus, the distribution of the target data (i.e., PLs) is fully determined by the teacher. The EMA update changes the PLs’ distribution in each iteration, resulting in unstable training. As a solution, we periodically update the teacher after a set number of iterations to maintain consistent distributions for PLs between updates. We call this strategy as the periodic update and show it outperforms the EMA update by a large margin in Sect. 4.3.
@@ 4 Experiments
ViLD [14] | 16+LSJ | RN50-FPN | FasterRCNN | 27.6 | 59.5 | 51.2 |
VL-PLM [57] | 1+Default | RN50-FPN | FasterRCNN | 32.3 | 54.0 | 48.3 |
F-VLM [24] | 0.5+LSJ | RN50-FPN | FasterRCNN | 28.0 | - | 39.6 |
@@
OVR-CNN [56] | (Not Given) | RN50-C4 | FasterRCNN | 22.8 | 46.0 | 39.9 |
RegionCLIP [58] | 1+Default | RN50-C4 | FasterRCNN | 26.8 | 54.8 | 47.5 |
Detic [59] | 1+Default | RN50-C4 | FasterRCNN | 27.8 | 51.1 | 45.0 |
PB-OVD [12] | 6+Default | RN50-C4 | FasterRCNN | 30.8 | 46.1 | 42.1 |
VLDet [29] | 1+LSJ | RN50-C4 | FasterRCNN | 32.0 | 50.6 | 45.8 |
BARON [48] | 1+Default | RN50-C4 | FasterRCNN | 33.1 | 54.8 | 49.1 |
SAS-Det (Ours) | 1+Default | RN50-C4 | FasterRCNN | 37.4 | 58.5 | 53.0 |
Method | Training Setup | Backbone | Detector | AP | AP | AP | AP |
ViLD [14] | 16+LSJ | RN50-FPN | FasterRCNN | 16.7 | 26.5 | 34.2 | 27.8 |
DetPro [10] | 2+Default | RN50-FPN | FasterRCNN | 20.8 | 27.8 | 32.4 | 28.4 |
F-VLM [24] | 9+LSJ | RN50-FPN | FasterRCNN | 18.6 | - | - | 24.2 |
BARON [48] | 2+Default | RN50-C4 | FasterRCNN | 17.3 | 25.6 | 31.0 | 26.3 |
SAS-Det (Ours) | 2+Default | RN50-C4 | FasterRCNN | 20.9 | 26.1 | 31.6 | 27.4 |
ViLD [14] | 16+LSJ | RN152-FPN | FasterRCNN | 19.8 | 27.1 | 34.5 | 28.7 |
@@
SAS-Det (Ours) | 2+Default | RN50x4-C4 | FasterRCNN | 29.1 | 32.4 | 36.8 | 33.5 |
@@ 4.1 Experiment Setup
Datasets. We conduct experiments on the two popular OVD benchmarks based on COCO [30] and LVIS [15]. For the COCO dataset, 65 out of 80 categories are divided into 48 base categories and 17 novel categories. We report results on the validation set. This setting is denoted as COCO-OVD. For the LVIS dataset, 337 rare categories are used as the novel concepts and the remaining 866 categories (frequent and common) as the base concepts. We report results on the whole LVIS validation set. This setting is denoted as LVIS-OVD.
Evaluation metrics. Following prior studies [56, 14, 58, 48], we report AP at IoU threshold of 0.5 for COCO-OVD. AP, AP and AP are the metrics for the novel, the base and all (novel + base) categories, respectively. For LVIS-OVD, we report the mean AP averaged on IoUs from 0.5 to 0.95. AP, AP, AP and AP are the metrics for rare, common, frequent, and all categories.
Implementation details. We built our detector on Faster RCNN with the CLIP version of ResNet (RN50-C4 or RN50x4-C4) as the backbone. Following prior studies [55, 57, 59], we assumed the novel concepts were available during training and took them as our open concepts, but no annotations of novel concepts are used. Our method was implemented on Detectron2 [50] and trained with 8 NVIDIA A6000 GPUs. We trained our models with the 1 schedule (90k iterations) for COCO-OVD and the 2 schedule for LVIS-OVD. The batch size was 16 with an initial learning rate of 0.002. The default data augmentation of Detectron2 was applied. Loss terms were equally weighted. By default, the teacher models were updated three times, usually together with the decreases of the learning rate.
@@ 4.2 Comparison with the Existing Methods
LVIS-OVD. We provide the main results on LVIS in Table 2. When using ResNet50 as the backbone, SAS-Det achieves similar performance as the recent method DetPro [10]. DetPro proposes learnable prompts to generate better text embeddings as the classifier for OVD. Our method is orthogonal and can leverage DetPro’s prompts for further improvement. The first block of Table 2 also shows that SAS-Det outperforms ViLD by a large margin on novel categories (indicated by AP) but gets lower AP for base categories. This is also observed in the recent method BARON [48]. The performance gap is probably due to the training setup. ViLD is trained for 8 times more iterations and adopts a strong data augmentation method [13], both of which benefit detection on base categories. In contrast, BARON and our approach adopt shorter training and standard data augmentations. When replacing the pretrained CLIP ResNet50 with ResNet50x4 [36] as the backbone, we improve AP by 8.2 from 20.9 to 29.1, which demonstrates SAS-Det scales up nicely with stronger pretrained VLM backbones.
@@ 4.3 Ablation Studies
Splitting the detection head. As shown in Table 3, compared to baseline, the open-branch of the proposed SAF head in (5) achieves similar performance on novel categories. This indicates that the split handles the noise in PLs’ location as well as the naive solution of baseline. This is plausible because both solutions exclude the noisy boxes of PLs from training.
Additionally, the open-branch gains 1.7 AP on base categories. Compared to (3)’s, the closed-branch in (4) gain improvements in terms of both AP and AP. Note that (4)’s and (3)’s heads are trained with the same data. Based on those results, the split benefits OVD beyond handling the location noise. One advantage is that, due to the split, the closed-branch is less likely to be influenced by the noise of PLs and learn how to better localize objects.
@@
The number of updates to the teacher model. We trained our detectors with different numbers of updates to teacher models. Generally, we consider the learning rate schedule and distribute updates as even as possible during training. Please see the supplement for more details about how we distribute the updates. As shown in Table 6, too many updates, e.g., 8 or 4 updates, lead to performance drops mainly due to the following. First, similar as the aforementioned EMA update, too many updates change the distribution of PLs too often and make the training unstable. Second, the more updates, the earlier an update happens. However, the student model is not well trained at the early stage of the training and thus is not good enough to update the teacher. Table 6 shows that 2 and 3 updates achieve similar performance. But we set 3 updates as default to include as many updates as possible, considering that the only overhead is to update the weights of the teacher with the student.
@@ 4.4 Further Analysis
Evaluation on the retained 15 COCO categories. Following prior studies [55, 57, 59], SAS-Det assumes novel concepts are available during training and achieves good performance. A natural question is if SAS-Det generalizes to alien concepts that are unknown before the evaluation. To answer the question, we follow the evaluation protocol of MEDet [7] where an OVD detector is evaluated on the whole COCO validation set with all 80 COCO concepts. AP, AP and AP are reported as APs averaged on 15 retrained, 17 novel and all 80 COCO categories, respectively. As shown in Table 6, SAS-Det achieves leading performance by a clear margin, which indicates the good generalization capability of the proposed method.
@@
Time cost of pseudo labeling. We run two major PL-based methods [12, 57] on our computational environment and compare them with our pseudo labeling in terms of the time cost in Table 3. As shown, our method is 3 faster than VL-PLM [57] and nearly 4 faster than PB-OVD [12], offering better PLs.
Reducing the time cost of external RPN. There is a simple solution to remove the extra computational cost of external RPN adopted by SAS-Det. We first generate pseudo labels with SAS-Det, and then follow VL-PLM [57] to train a Faster RCNN based on ResNet50 for OVD without the external RPN (“Faster RCNN + PLs”). As shown in Table 8, “Faster RCNN + PLs” achieves slightly better or close performance as SAS-Det.

### v3

@@ Abstract
Recent studies have shown promising performance in open-vocabulary object detection (OVD) by utilizing pseudo labels (PLs) from pretrained vision and language models (VLMs). However, teacher-student self-training, a powerful and widely used paradigm to leverage PLs, is rarely explored for OVD. This work identifies two challenges of using self-training in OVD: noisy PLs from VLMs and frequent distribution changes of PLs. To address these challenges, we propose SAS-Det that tames self-training for OVD from two key perspectives. First, we present a split-and-fusion (SAF) head that splits a standard detection into an open-branch and a closed-branch. This design can reduce noisy supervision from pseudo boxes. Moreover, the two branches learn complementary knowledge from different training data, significantly enhancing performance when fused together. Second, in our view, unlike in closed-set tasks, the PL distributions in OVD are solely determined by the teacher model. We introduce a periodic update strategy to decrease the number of updates to the teacher, thereby decreasing the frequency of changes in PL distributions, which stabilizes the training process. Extensive experiments demonstrate SAS-Det is both efficient and effective. SAS-Det outperforms recent models of the same scale by a clear margin and achieves 37.4 AP50 and 29.1 APr on novel categories of the COCO and LVIS benchmarks, respectively. Code is available at https://github.com/xiaofeng94/SAS-Det.
@@ 1 Introduction
Traditional closed-set object detectors [6, 16, 40] are restricted to detecting objects with a limited number of categories. Increasing the size of detection vocabularies usually requires heavy human labor to collect annotated data. With the recent advent of strong vision and language models (VLMs) [26, 38], open-vocabulary object detection (OVD) [14] provides an alternative direction to approach this challenge. Typically, OVD detectors are trained with annotations of base categories and expected to generalize to novel categories with the power of pretrained VLMs.
One promising thread of recent studies for OVD [11, 12, 59, 51] leverages VLMs to obtain pseudo labels (PLs) beyond base categories. But they rarely explore self-training, a powerful and widely used schema for utilizing PLs in closed-set tasks [47, 46, 53, 54]. We investigate this and find the vanilla self-training approach does not improve OVD performance due to the following challenges.
@@
Second, self-training for closed-set object detection [47, 54, 25] usually follows an online teacher-student manner. In each training iteration, the teacher generates PLs, and the student is trained with a mixture of ground truth and PLs. Then, the teacher is updated by the student with exponential moving average (EMA). However, we find such EMA updates degrade OVD models (see Table 4). Our hypothesis is that, unlike closed-set tasks, OVD provides no ground truth for target categories, and thus, the supervision for target categories is fully decided by the distribution of PLs predicted by the teacher. Hence, the EMA updates change the distribution of PLs in each iteration, unstabilizing the training.
@@ 2 Related Work
Open-vocabulary object detection (OVD). Zero-shot object detection methods [3, 39, 62, 55] increase the vocabulary size but with limited accuracy. Motivated by the strong zero-shot abilities of VLMs, recent efforts focus on OVD. Finetuning-based methods [58, 34, 60, 24, 22] add detection heads onto pretrained VLMs and then finetune the detector with concepts of base categories. Such methods are simple but may forget the knowledge learned in the pretraining [2]. Distillation-based methods [14, 10, 50, 49] introduce additional distillation loss functions that force the output of a detector to be close to that of a VLM and thus avoid forgetting. However, since the distillation losses are not designed for the detection task, they may conflict with detection objectives due to gradient conflicts that are a common issue for multi-task models. Methods like Gao et al. [12], Zhao et al. [59], Feng et al. [11], Wu et al. [51] create pseudo labels (PL) of novel concepts as supervision, and do not require extra losses, which sidesteps both catastrophic forgetting and gradient conflicts. But these methods need handcrafted steps to generate high quality PLs, e.g. multiple runs of box regression [59], activation maps [12] from Grad-CAM [43], image retrieval [11], or multi-stage training [51]. In this work, we address the two challenges of using self-training for OVD and enable an efficient end-to-end pseudo labeling pipeline. Besides, we point out the noise due to the poor locations of PLs and introduce the SAF head to handle such noise.
Self-training for object detection. Weakly-supervised object detection methods [47, 48, 41, 33] usually explore online self-training by distilling the knowledge from the model itself. Recent semi-supervised object detection methods [54, 32, 25] adopt a teacher-student design, where the teacher is an exponential moving average (EMA) of the student. Self-training has been widely explored in the above fields but rarely in OVD. Unlike semi-supervised object detection, OVD encounters two challenges to use self-training, i.e., more noisy PLs, and larger changes in PLs’ distributions. This work proposes the SAF head to address the noise, and adopts periodic updates to reduce the frequency of changes in PLs’ distribution.
@@ 3.1 Adapting CLIP to OVD
In this section, we introduce how to adapt the ResNet-based CLIP into a Faster R-CNN [40] (C4) detector. For simplicity, we do not use FPN [31], but it can be incorporated with learnable gating from Flamingo [1]. Similarly, ViT-based CLIP models can be adapted to detectors following [28].
@@ 3.2 Taming Self-Training
Reducing changes in PLs’ distributions with periodic updates. The teacher model is updated with the student to enable self-training. The exponential moving average (EMA) is a widely used approach for object detection [25, 32, 54], which updates the teacher in every iteration. But we observed empirically that it does not benefit OVD (Table 4). We hypothesize that, unlike semi-supervised object detection, OVD has no ground truth for the target categories. Thus, the distribution of the target data (i.e., PLs) is fully determined by the teacher. The EMA update changes the PLs’ distribution in each iteration and leads to a constantly shifting training target that has been shown hard to optimize (i.e., destabilizing the training) in deep Q-learning [35]. As a solution, we periodically update the teacher after a set number of iterations to maintain consistent distributions for PLs between updates. We call this strategy as the periodic update and show it outperforms the EMA update by a large margin in Sect. 4.3.
@@ 4 Experiments
ViLD [14] | 16+LSJ | RN50-FPN | Faster R-CNN | 27.6 | 59.5 | 51.2 |
VL-PLM [59] | 1+Default | RN50-FPN | Faster R-CNN | 32.3 | 54.0 | 48.3 |
F-VLM [24] | 0.5+LSJ | RN50-FPN | Faster R-CNN | 28.0 | - | 39.6 |
@@
OVR-CNN [58] | (Not Given) | RN50-C4 | Faster R-CNN | 22.8 | 46.0 | 39.9 |
RegionCLIP [60] | 1+Default | RN50-C4 | Faster R-CNN | 26.8 | 54.8 | 47.5 |
Detic [61] | 1+Default | RN50-C4 | Faster R-CNN | 27.8 | 51.1 | 45.0 |
PB-OVD [12] | 6+Default | RN50-C4 | Faster R-CNN | 30.8 | 46.1 | 42.1 |
VLDet [29] | 1+LSJ | RN50-C4 | Faster R-CNN | 32.0 | 50.6 | 45.8 |
BARON [50] | 1+Default | RN50-C4 | Faster R-CNN | 33.1 | 54.8 | 49.1 |
SAS-Det (Ours) | 1+Default | RN50-C4 | Faster R-CNN | 37.4 | 58.5 | 53.0 |
Method | Training Setup | Backbone | Detector | APr | APc | APf | AP |
ViLD [14] | 16+LSJ | RN50-FPN | Faster R-CNN | 16.7 | 26.5 | 34.2 | 27.8 |
DetPro [10] | 2+Default | RN50-FPN | Faster R-CNN | 20.8 | 27.8 | 32.4 | 28.4 |
F-VLM [24] | 9+LSJ | RN50-FPN | Faster R-CNN | 18.6 | - | - | 24.2 |
BARON [50] | 2+Default | RN50-C4 | Faster R-CNN | 17.3 | 25.6 | 31.0 | 26.3 |
SAS-Det (Ours) | 2+Default | RN50-C4 | Faster R-CNN | 20.9 | 26.1 | 31.6 | 27.4 |
ViLD [14] | 16+LSJ | RN152-FPN | Faster R-CNN | 19.8 | 27.1 | 34.5 | 28.7 |
@@
SAS-Det (Ours) | 2+Default | RN50x4-C4 | Faster R-CNN | 29.1 | 32.4 | 36.8 | 33.5 |
@@ 4.1 Experiment Setup
Datasets. We conduct experiments on the two popular OVD benchmarks COCO-OVD and LVIS-OVD, based on COCO [30] and LVIS [15], respectively. (1) COCO-OVD: 65 out of 80 COCO categories are divided into 48 base categories and 17 novel categories. We report results on the validation set. (2) LVIS-OVD: 337 rare categories are used as the novel concepts and the remaining 866 categories (frequent and common) as the base concepts. We report results on the whole LVIS validation set.
Evaluation metrics. Following prior studies [58, 14, 60, 50], we report AP at IoU threshold of 0.5 for COCO-OVD. AP, AP and AP are the metrics for the novel, the base and all (novel + base) categories, respectively. For LVIS-OVD, we report the mean AP averaged on IoUs from 0.5 to 0.95. APr, APc, APf and AP are the metrics for rare, common, frequent, and all categories.
Implementation details. We built our detector on Faster R-CNN with the CLIP version of ResNet (RN50-C4 or RN50x4-C4) as the backbone. Following prior studies [57, 59, 61], we assumed the novel concepts were available during training and took them as our open concepts, but no annotations of novel concepts are used. Our method was implemented on Detectron2 [52] and trained with 8 NVIDIA A6000 GPUs. We trained our models with the 1 schedule (90k iterations) for COCO-OVD and the 2 schedule for LVIS-OVD. The batch size was 16 with an initial learning rate of 0.002. The default data augmentation of Detectron2 was applied. Loss terms were equally weighted. By default, the teacher models were updated three times, usually together with the decreases of the learning rate.
@@ 4.2 Comparison with the Existing Methods
LVIS-OVD. We provide the main results on LVIS in Table 2. When using ResNet50 as the backbone, SAS-Det achieves similar performance as the recent method DetPro [10]. DetPro proposes learnable prompts to generate better text embeddings as the classifier for OVD. Our method is orthogonal and can leverage DetPro’s prompts for further improvement. The first block of Table 2 also shows that SAS-Det outperforms ViLD by a large margin on novel categories (indicated by APr) but gets lower APf for base categories. This is also observed in the recent method BARON [50]. The performance gap is probably due to the training setup. ViLD is trained for 8 times more iterations and adopts a strong data augmentation method [13], both of which benefit detection on base categories. In contrast, BARON and our approach adopt shorter training and standard data augmentations. When replacing the pretrained CLIP ResNet50 with ResNet50x4 [38] as the backbone, we improve APr by 8.2 from 20.9 to 29.1, which demonstrates SAS-Det scales up nicely with stronger pretrained VLM backbones.
@@ 4.3 Ablation Studies
Splitting the detection head. As shown in Table 3, compared to baseline, the open-branch of the proposed SAF head in (5) achieves similar performance on novel categories. This indicates that the split handles the noise in PLs’ location as well as the naive solution of baseline. This is plausible because both solutions exclude the noisy boxes of PLs from training. Additionally, the open-branch gains 1.7 AP on base categories. Compared to (3)’s, the closed-branch in (4) gain improvements in terms of both AP and AP. Note that (4)’s and (3)’s heads are trained with the same data. Based on those results, the split benefits OVD beyond handling the location noise. One advantage is that, due to the split, the closed-branch is less likely to be influenced by the noise of PLs and learn how to better localize objects.
@@
The number of updates to the teacher model. We trained our detectors with different numbers of updates to teacher models. Please see the supplement for how we distribute the updates. As shown in Table 6, too many updates, e.g., 8 or 4 updates, lead to performance drops mainly due to the following. First, similar as the aforementioned EMA update, too many updates change the distribution of PLs too often and make the training unstable. Second, the more updates, the earlier an update happens. However, the student model is not well trained at the early stage of the training and thus is not good enough to update the teacher. Table 6 shows that 2 and 3 updates achieve similar performance. But we set 3 updates as default to include as many updates as possible.
@@ 4.4 Further Analysis
Evaluation on the retained 15 COCO categories. Following prior studies [57, 59, 61], SAS-Det assumes novel concepts are available during training and achieves good performance. A natural question is if SAS-Det generalizes to alien concepts that are unknown before the evaluation. To answer the question, we follow the evaluation protocol of MEDet [7] where an OVD detector is evaluated on the whole COCO validation set with all 80 COCO concepts. AP, AP and AP are reported as APs averaged on 15 retained, 17 novel, and all 80 COCO categories, respectively. As shown in Table 6, SAS-Det achieves leading performance by a clear margin, which indicates the good generalization capability of the proposed method.
@@
Time cost of pseudo labeling. We run two major PL-based methods [12, 59] on our computational environment and compare them with our pseudo labeling in terms of the time cost in Table 7.
Reducing the time cost of external RPN. There is a simple solution to remove the extra computational cost of external RPN adopted by SAS-Det. We first generate pseudo labels with SAS-Det, and then follow VL-PLM [59] to train a Faster R-CNN based on ResNet50 for OVD without the external RPN (“Faster R-CNN + PLs”). As shown in Table 8, “Faster R-CNN + PLs” achieves slightly better or close performance as SAS-Det.